In 7 year asim will be twice as old as he was 8 year ago.how old he is now?
Found 2 solutions by macston, fractalier:
Answer by macston(5194)

.
A=Asim's age
.
A+7=2(A-8)
A+7=2A-16
23=A
.
ANSWER: Asim is 23 years old.
.

Answer by fractalier(6550)

Call his age NOW, x. Then we have
x + 7 = 2(x - 8)
x + 7 = 2x - 16
x = 23